OpenStudy (anonymous):

Write the equation using function notation through (3,8) parallel to f(x)=4x-2

OpenStudy (alexandervonhumboldt2):

so parralel lines have same slope what is slope for the line parallel to f(X)?

OpenStudy (anonymous):

4

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (alexandervonhumboldt2):

so write y=4x+b

OpenStudy (alexandervonhumboldt2):

to find y-intersept substitue point: 8=4*3+b b=8-12=?

OpenStudy (anonymous):

-4

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (alexandervonhumboldt2):

so the equation is y=4x-4

OpenStudy (anonymous):

Great, thank you!

OpenStudy (alexandervonhumboldt2):

np thz for medal can you also fan plz?